Number plates in an Arab country consist of 3 letters (chosen from a specified set of 20 letters from the Arabic alphabet) and 4 digits (from 0 to 9).


Find the probability of a plate having identical letters.


Find the probability of a plate having identical digits.


1
Expert's answer
2022-04-14T17:18:30-0400

"1:\\20\cdot \frac{1}{20}\cdot \frac{1}{20}\cdot \frac{1}{20}=0.0025\\2:\\10\cdot \frac{1}{10}\cdot \frac{1}{10}\cdot \frac{1}{10}\cdot \frac{1}{10}=0.001\\"